OpenStudy (anonymous):

Geometry help!! I'm stuck In triangle ABC, Angle A=2x+3, Angle B=3x+5, and Angle C=5x+2. Solve for x and find the measure of each angle. I don't know what to do here... do I make it 2x+3+3x+5+5x+2=180?...

OpenStudy (anonymous):

or do I do 1 at a time 2x+3=180 5x+2=180 and 3x+5=180..?